### v2

@@ 1 Background
Minimally invasive surgery (MIS), including laparoscopic and robot-assisted procedures, has transformed surgical practice by reducing patient recovery time, minimizing tissue trauma, and improving cosmetic outcomes (Mack, 2001). As these techniques become the standard of care in a wide range of clinical specialties, from general surgery to gynecology and urology, the demand for enhanced surgical vision and intra-operative guidance has increased in parallel (Conrad et al., 2015; Oberlin et al., 2016). In particular, the quality of intra-operative imaging has become increasingly vital for ensuring surgical safety, precise dissection, and the detection of critical anatomical features such as vessel bifurcations, lesion margins, or hidden pathology (Marshall et al., 2010; Goh et al., 2010).
@@
In recent years, advances in medical imaging have driven a shift toward higher-resolution acquisition systems. In the domain of MIS, modern laparoscopy platforms such as Olympus VISERA or Stryker 1688 have begun offering ultra-high-definition (UHD) 4K images, promising substantial benefits in surgical navigation and decision making (Abdelrahman et al., 2018). High resolution visualization enables surgeons to better differentiate tissue types, observe microvasculature, and assess resection margins, capabilities that are crucial for both oncologic and reconstructive interventions, as shown in Figure 1. Recent clinical and simulator studies demonstrate that 4k laparoscopy shortens the operative time and reduces blood loss compared to HD systems (Puccetti et al., 2023). Randomized skill acquisition trials also show faster learning curves and fewer errors for 4K vision (Abdelrahman et al., 2018), and systematic evidence confirms that many performance metrics are close to those of dedicated 3D systems when 4K is used (Harada et al., 2018). For example, accurate margin assessment during colorectal or liver surgery can directly impact long-term patient outcomes (Montalti et al., 2015). In addition, detailed imaging facilitates the development of downstream computer-assisted intervention (CAI) tools, such as automated segmentation, tissue classification, and image-based depth estimation. 4K resolution is widely available in non-surgical commercial space, yet most public surgical vision instruments remain trapped in the HD/sub-HD era, preventing algorithms from unleashing their full potential on utilizing high-resolution images.
Figure 2: Example frames from the training dataset, showcasing various tools and scenarios used in different scenarios. These frames highlight the diversity of situations included in the dataset.
@@ 3.2 Data Location
The raw SurgiSR4K video clips and associated high-resolution images contain proprietary and sensitive surgical content and are therefore available only upon request. Researchers may request access by email with a brief description of their intended use. Requests will be reviewed, and, if approved, a data-sharing agreement will be provided. All derived datasets and the complete code base will be publicly available.
@@ 3.4 Ethical Considerations
This is an animal model data set collected in vivo from healthy Yorkshire swines (weight 25 to 35 kg) under general anesthesia. No human subjects were involved; demographic data is not applicable. Each session corresponds to a distinct surgical procedure with fresh bowel preparation. Inclusion criterion: normal gastrointestinal tract; exclusion: previous surgical alteration or pathology. The procedures were performed under the supervision of expert veterinarians.

### v3

@@ 1 Background
Minimally invasive surgery (MIS), including laparoscopic and robot-assisted procedures, has transformed surgical practice by reducing patient recovery time, minimizing tissue trauma, and improving cosmetic outcomes (Mack, 2001). As these techniques become the standard of care in a wide range of clinical specialties, from general surgery to gynecology and urology, the demand for enhanced surgical vision and intra-operative guidance has increased in parallel (Conrad et al., 2015; Oberlin et al., 2016). In particular, the quality of intra-operative imaging has become increasingly vital for ensuring surgical safety, precise dissection, and the detection of critical anatomical features such as vessel bifurcations, lesion margins, or hidden pathology (Marshall et al., 2010; Goh et al., 2010). For example, accurate margin assessment during colorectal or liver surgery can directly impact long-term patient outcomes (Montalti et al., 2015).
@@
In recent years, advances in medical imaging have driven a shift toward higher-resolution acquisition systems. In the domain of MIS, modern laparoscopy platforms such as Olympus VISERA or Stryker 1688 have begun offering ultra-high-definition (UHD) 4K images, promising substantial benefits in surgical navigation and decision making (Abdelrahman et al., 2018). High resolution visualization enables surgeons to better differentiate tissue types, observe microvasculature, and assess resection margins, capabilities that are crucial for both oncologic and reconstructive interventions, as shown in Figure 1. Recent clinical and simulator studies demonstrate that 4k laparoscopy shortens the operative time and reduces blood loss compared to HD systems (Puccetti et al., 2023). Randomized skill acquisition trials also show faster learning curves and fewer errors for 4K vision (Abdelrahman et al., 2018), and systematic evidence confirms that many performance metrics are close to those of dedicated 3D systems when 4K is used (Harada et al., 2018). In addition, detailed imaging facilitates the development of downstream computer-assisted intervention (CAI) tools, such as automated segmentation, tissue classification, and image-based depth estimation. 4K resolution is widely available in non-surgical commercial space, yet most public surgical vision instruments remain trapped in the HD/sub-HD era, preventing algorithms from unleashing their full potential on utilizing high-resolution images.
Figure 2: Example frames from the training dataset, showcasing various tools used in different scenarios. These frames highlight the diversity of situations included in the dataset.
@@ 3.2 Data Location
The SurgiSR4K dataset—including video clips and corresponding high-resolution images—is publicly available through our dedicated repository on Synapse:syn68756003. The portal provides open access to the complete dataset, along with detailed metadata, project wiki, and supporting documentation.
@@ 3.4 Ethical Considerations
This animal model dataset was collected in vivo from healthy Yorkshire swine under general anesthesia, with no human subjects involved.
All procedures were approved and conducted in accordance with Intuitive Surgical’s Institutional Animal Care and Use Committee (IACUC) protocol No. IACUC‑2022‑1200 (Title: [004] Use of Porcine Models for Robotic Surgical and Procedure Training). The study was conducted in alignment with the 3Rs principles—Refinement, Reduction, and Replacement—to ensure the ethical and humane treatment of animals.
All procedures were supervised by qualified veterinary staff, and personnel involved were trained and approved under the relevant IACUC protocols.